Suppose a car jack has an input work 2750 J. It lifts a car of mass 750 kg a distance of 0.35 meters in 8 seconds. What is the output work (Wout) of the jack? How much power is required to lift the car? What is the efficiency of the jack?
Answer #1

Work done by jack =Wout = Force * displacement = mg * h = 750 * 9.8 * 0.35 = 2572.5 J

Power required = work done/ time taken = 2572.5/8 = 321.56 W

Efficiency = (energy for work) /(input energy) = 2572.5/2750 = 0.9354 = 93.54%